I'm having trouble solving the following:
7x+3(5-2x)=21
I know x=6 but I'm having difficulty figuring out the steps.
Can you start me out in the right direction?
I don't know how to get rid of the x in the parenthesis.
Any help you could give me would be appreciated.

Answer by Lal(16)

7x + 3(5 - 2x) = 21

For the parentheses, you have to distribute the 3 to the numbers inside of them. So, it would look like this, after distribution:

7x + 15 - 6x = 21

And then, to solve:

7x - 6x = 6
1x = 6
x = 6